OpenStudy (anonymous):

y = -cot(x + pie/6) - 3 graphing a trig function please help me!!!!!

OpenStudy (anonymous):

thx its just a little confusing on there

OpenStudy (anonymous):

take the graph of \(y=\cot(x)\) shift it to the left \(\frac{\pi}{6}\) units that gives \(y=\cot(x+\frac{\pi}{6})\) then turn it upside down to get \(y=-\cot(x+\frac{\pi}{6})\) then translate it down 3 units to get \[y=-\cot(x+\frac{\pi}{6})-3\]

OpenStudy (anonymous):

one question tho

OpenStudy (anonymous):

like if im starting at pie/4 interval how can i get to pie/6 ?

OpenStudy (anonymous):

you can't

OpenStudy (anonymous):

so than like how do i lable the graph its so confusing

OpenStudy (anonymous):

there is no \(\frac{\pi}{4}\) in the problem

OpenStudy (anonymous):

the interval

OpenStudy (anonymous):

here is a picture of \(y=\cot(x)\) you see it has a vertical asymptote at \(0\) and at \(\pi\) and also that it is zero at \(\frac{\pi}{2}\)

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=cot%28x%29

OpenStudy (anonymous):

you want to shift it to the left \(\frac{\pi}{6}\) units, so instead of the vertical asymptote at \(0\) and \(\pi\) they will be at \(-\frac{\pi}{6}\) and \(\frac{5\pi}{6}\)

OpenStudy (anonymous):

-cot and cot are the same thing?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i was just trying to get the graph of \(y=\cot(x+\frac{\pi}{6})\) first , then we can turn it upside down

OpenStudy (anonymous):

i.e. see that you are going to shift it to the left \(\frac{\pi}{6}\) units here it is http://www.wolframalpha.com/input/?i=cot%28x%2Bpi%2F6%29

OpenStudy (anonymous):

you can compare the graphs and see that the second one is identical to the first, only shifted left \(\frac{\pi}{6}\) now we can flip it

OpenStudy (anonymous):

last, question how would u label it and thanks a lot

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=-cot%28x%2Bpi%2F6%29 asymptotes are the same, now it is upside down

OpenStudy (anonymous):

and finally we can drop it by 3 units to finish the question, and get this graph http://www.wolframalpha.com/input/?i=-cot%28x%2Bpi%2F6%29-3

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

if you want to label the \(x\) axis, you probably best label the zero and the asymptotes

OpenStudy (anonymous):

yea im just confused, if it starts at -pi/6 do i label it by going up pi/6?

OpenStudy (anonymous):

instead of it being at \(0\) and \(\pi\) as i said before, label the asymptotes at \(-\frac{\pi}{6}\) and \(\frac{5\pi}{6}\)

OpenStudy (anonymous):

ok yea

OpenStudy (anonymous):

and 11pi/6

OpenStudy (anonymous):



OpenStudy (anonymous):

right, the next one would be \(\frac{11\pi}{6}\)

OpenStudy (anonymous):